# FAST: An Efficient Scheduler for All-to-All GPU Communication

### v2

@@ 1 Introduction
All-to-All communication—where every endpoint sends data to all others—has long been a fundamental collective communication primitive in scientific and parallel computing, supporting workloads such as 3D FFTs [41]. Early systems typically treated each server as the communication endpoint, interconnected by networks with relatively uniform bandwidth. In modern ML clusters, the endpoint has shifted to individual GPUs, which are connected through a two-tier fabric consisting of faster intra-server (scale-up) links and slower inter-server (scale-out) links.111In this paper, the terms ‘scale-up’ and ‘intra-server’ network are used interchangeably, and likewise ‘scale-out’ and ‘inter-server.’ Consequently, All-to-All has become a key operation for many ML applications, including recommendation models [31, 32], Gaussian Splatting [47], and mixture-of-experts (MoE) models [43, 23, 14, 20].
@@
The role of All-to-All in ML applications is important. In Mixture-of-Experts (MoE) models in particular, its cost can account for a large fraction of training time. MoE improves the efficiency of model parameters by activating only a subset of experts for each input token rather than all simultaneously. This selectivity, however, necessitates frequent All-to-All operations to dispatch tokens to experts across GPUs and aggregate their outputs. Prior studies [25, 20] show that MoE All-to-All can consume 30–55% of training time, making it a major contributor to overhead in large-scale training.
@@
The challenges of All-to-All arise at both the application and system layers. At the application layer, traffic is often skewed and dynamic. In MoE, some experts are selected more frequently than others, leading to larger data transfers for their corresponding GPUs during All-to-All. This imbalance keeps certain GPUs and NICs busy long after others have finished, creating straggler effects. When communication volumes differ across endpoints in All-to-All, the operation is referred to as alltoallv [30]. Further, the traffic pattern in alltoallv changes every few hundred milliseconds, as the MoE gating function reassigns tokens to experts at runtime (Figure˜1). As a result, a GPU that is a hotspot at one moment may be nearly idle the next. This dynamism makes static alltoallv schedules impractical and requires schedulers to adapt in real time.
@@
At the system layer, the hardware fabric that connects GPUs is inherently two-tiered: fast intra-server links (scale-up) and much slower inter-server links (scale-out) (Figure˜4). This heterogeneity means that flows of the same size can finish quickly inside a server but take an order of magnitude longer across servers, leaving schedulers to coordinate thousands of flows over mismatched bandwidths. In addition, alltoallv’s dense communication pattern naturally triggers incast, a classic networking problem where many senders overload downlink of the same receiver. Incast causes network congestion with switch queue buildup and reduced goodput—even under modern congestion control—and remains an open challenge in large-scale cluster networking [12, 19].
@@
Together, these challenges make efficient alltoallv scheduling especially difficult under the tight timescales demanded by MoE. Existing schedulers such as TACCL [44] and TE-CCL [27] employ solvers [18] to generate near-optimal schedules. While they overcome the inefficiencies of fixed schedules in collective communication libraries such as NCCL and RCCL, their resulting formulations are NP-hard [44] and can take minutes to hours to synthesize a schedule even for just 32 GPUs. The state-of-the-art scheduler SyCCL [10] accelerates this process with heuristics and parallelism, yet require seconds to minutes for balanced All-to-All, while skewed alltoallv still remains unresolved. Consequently, these schedulers are far too slow for MoE alltoallv workloads that shift every few hundred milliseconds.
@@
In this paper, rather than increasing scheduler complexity, we simplify the problem itself: It suffices to focus on optimizing the scale-out tier—the real bottleneck. We observe that scale-up is roughly an order-of-magnitude faster than scale-out (Footnote˜3). So the much faster scale-up fabric can cheaply absorb skew within each server, reshaping traffic before it reaches scale-out. With this, we can then retain best possible performance by keeping bottleneck servers transmitting at full rate while avoiding incast—maximizing scale-out efficiency. This naturally reduces to a one-to-one matching problem between endpoints, solvable with polynomial-time algorithms.
@@
Building on this insight, we propose FAST, a polynomial-time, matching-based scheduler for skewed and dynamic alltoallv workloads. FAST operates in two phases: (i) Skew mitigation, where it uses fast scale-up fabric to rebalance the scale-out workload so that all NICs face equal volume before traversing the slow scale-out links; and (ii) Balanced, one-to-one transfers, which use Birkhoff’s decomposition [8] to generate successive matchings between senders and receivers, ensuring scale-out transfers proceed without incast while keeping bottleneck servers fully active at line rate until completion. While Birkhoff’s decomposition has appeared in the design of crossbar switches [26, 11], this is, to our knowledge, the first application to collective communication.
@@
We implement FAST on both NVIDIA H200 [38] and AMD MI300X [2] testbeds and evaluate it against state-of-the-art solutions such as DeepEP [46], TACCL [44], and TE-CCL [27]. Under skewed workloads, FAST outperforms the strongest NVIDIA baseline by 1.01–1.3 and the strongest AMD baselines by 1.5–2.8, and when integrated into Megatron-LM [45] on AMD, improves end-to-end MoE training throughput by over RCCL [5] that suffers heavily from incast. The scheduler is highly efficient, completing in 25 s for 32 GPUs, 221 s for 64 GPUs—fast enough for MoE workloads where the traffic matrix changes every few hundred milliseconds, whereas solver-based methods require minutes to hours even at 32 GPUs. Our code will be made publicly available.
@@ 2 Motivation
All-to-All (alltoallv) communication—where every endpoint exchanges distinct data with all others—has become a key primitive in modern GPU clusters. In this setting, the communication endpoint is the individual GPU, connected by a two-tier fabric consisting of fast intra-server (scale-up) links and slower inter-server (scale-out) links.
@@
All-to-All communication cost in MoE models. Mixture-of-experts (MoE) is now a leading architecture for scaling large language models: instead of activating the full model for every input, a lightweight gating network selects only a subset of ‘experts.’ This improves model parameter efficiency but requires frequent, large-scale alltoallv operations to dispatch tokens to their chosen experts and then gather results. As illustrated in Figure˜1, each MoE layer invokes alltoallv twice, and MoE layers often constitute a large fraction of the model. Prior measurements [20, 25] have shown that alltoallv can account for 30–56% of training time.
@@
While MoE is our primary focus, the importance of alltoallv extends beyond. It underpins recommendation systems [31, 32], Gaussian Splatting [47], and classical scientific workloads such as 3D FFT(where it can dominate up to 97.3% of runtime [7]!). Despite being just one collective, alltoallv performance disproportionately shapes the efficiency of both modern AI and traditional HPC workloads.
@@
Application challenges—skewness & dynamism. To study this communication, we profile MoE training using Megatron-LM [45] with 32 experts (one per GPU). We find that alltoallv workloads are inherently skewed and dynamic, consistent with recent profiling results of Mixtral models [25, 21]. Unlike balanced collectives such as All-Reduce, MoE alltoallv generates a highly uneven demand matrix: some GPU pairs exchange more than 12 the median volume (Figure˜2(a)). This skew creates stragglers—NICs that remain busy long after others have finished, delaying the entire collective and stalling training progress (Figure˜3(a)). In this setting, scheduling is critical: by routing part of a heavy GPU’s traffic to idle NICs, a smart scheduler can smooth out skew and mitigate stragglers, as illustrated in Figure˜3(b).
@@
The MoE workload is also inherently dynamic: the alltoallv traffic pattern shifts every few hundred milliseconds. As illustrated in Figure˜2(b), a GPU that is overloaded in one alltoallv may become nearly idle in the next, since token routing is determined jointly by the input tokens and the per-MoE-layer gating functions (Figure˜1). Because these patterns cannot be predicted in advance, schedules must be recomputed online at the same frequency as workload changes—making fast, on-the-fly scheduling essential.
@@
System challenges—heterogeneity & incast. Modern GPU clusters further complicate scheduling with two system-level challenges. First, a heterogeneous, two-tier fabric connects GPUs (Figure˜4(a)): fast intra-server links (e.g., NVLink, up to 900 GBps) and much slower inter-server links (e.g., InfiniBand, 400–800 Gbps). This two-tier network increases scheduling complexity: for each alltoallv, schedulers must navigate thousands of flows across mismatched bandwidths, exploring a large space of routing and pacing decisions, which can quickly turn scheduling itself into a bottleneck.
@@
Second, incast is a classic networking problem that arises from alltoallv’s dense communication pattern, where many flows converge on the same NIC downlink in the scale-out fabric. While the burstiness of small messages can be absorbed by switch queues, MoE alltoallv transfers are much larger—typically 100 MB to 1 GB [46]—causing sustained congestion that requires active control. Even with advanced schemes [17, 29, 19], incast often results in unfair bandwidth sharing and degraded goodput. It remains an open challenge [12], and schedulers often attempt to mitigate it proactively rather than relying on the transport layer alone.
@@
Limitations of existing approaches. State-of-the-art schedulers such as TACCL [44], TE-CCL [27], and SyCCL [10] are designed to be general: they support a wide range of collectives (All-Reduce, All-Gather, All-to-All, etc.) and reason about arbitrary topologies. To achieve this generality, they cast scheduling as constraint-satisfaction or optimization problems, often NP-hard [44], and solve them with heavy computation. This is practical for collectives with repetitive communication patterns like All-Reduce, where the high scheduling cost can be amortized over many iterations.
@@
For alltoallv, however, existing approaches are far too slow. SyCCL [10], the fastest to date with parallelism and heuristic acceleration, is orders of magnitude faster than earlier solver-based systems but still requires minutes to produce a schedule for 64 GPUs—impractical when MoE traffic shifts every few hundred milliseconds. While such systems excel at achieving near-optimal completion under arbitrary topologies, their fine-grained modeling makes them hardly scalable to this setting.
@@
At the other extreme, production libraries like NCCL [3] and MPI [33] generate schedules instantly, but rely on fixed schedules oblivious to the dynamic, skew workload, often resulting in much lower throughput than what the hardware could achieve.
@@
Goal. Can we design a fast, online scheduler for dynamic, skewed alltoallv that mitigates stragglers and sustains high performance? Rather than targeting all collectives or arbitrary topologies, our goal is a specialized solution for alltoallv on today’s two-tier GPU clusters—where existing schedulers struggle to handle the underlying skewness, dynamism, asymmetry, and incast challenges.
@@ 3 Design Overview
To build a fast, on-the-fly scheduler for alltoallv in modern GPU clusters, we take a step back. Rather than enumerating complex constraints, we start by first solving alltoallv on a simplified single-tier network and then generalize the solution step by step to today’s asymmetric two-tier fabrics.
Starting point: alltoallv in a single-tier network. On a single-tier, full-bisection network where all links have identical bandwidth, the goal is clear: avoid incast and congestion by ensuring that, at any given time, each sender communicates with exactly one receiver, and each receiver accepts data from exactly one sender. This naturally leads to a staged schedule: in each stage, senders and receivers are matched one-to-one, and across stages, all data is exchanged across every sender–receiver combination.
@@
To solve this problem, we observe that Birkhoff’s decomposition [8]—introduced in 1946 as a purely mathematical identity—can be reinterpreted as an optimal scheduling strategy for alltoallv. Formally, the theorem states that any traffic matrix444The theorem applies to scaled doubly stochastic matrices; arbitrary matrices can be adapted, as described later. can be expressed as a weighted sum of permutation matrices. Viewed through the scheduling lens, each permutation corresponds to a transfer stage: every active row (sender) and column (receiver) has exactly one nonzero entry of equal size (transfer size), so each participant exchanges data with exactly one partner and all finish the stage together. By summing over these permutation matrices, all flows advance in proportion to their demand, driving the transfer to completion.
Figure˜5 illustrates our strawman aproach: the top pane shows the decomposition into (partial) permutation matrices, while the bottom pane shows the corresponding transfer schedule. This approach is appealing because (i) completion time matches the lower bound—the bottleneck node (e.g., as sender in blue) transmits in all stages; (ii) each stage is balanced until nodes finish, so participants advance to the next stage together; and (iii) the decomposition is computationally efficient.
Challenges of applying Birkhoff’s decomposition to heterogeneous two-tier networks. Unfortunately, today’s GPU clusters deviate from the simplified network setting in two ways: (i) In a two-tier fabric, even a ‘balanced’ permutation stage can finish unevenly: transfers on the slow scale-out links lag behind, idling the faster scale-up links and wasting bandwidth; (ii) More critically, completion time is dictated by the busiest scale-out NIC: the heaviest sender or receiver. If skew is large, Birkhoff’s schedule still stalls at those endpoints.555Intra-server traffic rarely dominates in multi-node ML workloads, since the number of scale-out pairs is far larger than scale-up pairs.
@@
Inter-server scheduling: balanced one-to-one transfers (§4.2). Once intra-server skew is absorbed, the remaining challenges are server-level imbalance and incast. Here, we apply Birkhoff’s decomposition to construct successive one-to-one, balanced server-to-server transfer stages, ensuring bottleneck servers remain continuously active at line rate until their traffic is complete.

### v3

@@ 1 Introduction
All-to-All communication—where every endpoint sends data to all others—has long been a fundamental collective communication primitive in scientific and parallel computing, supporting workloads such as 3D FFTs [fft]. Early systems typically treated each server as the communication endpoint, interconnected by networks with relatively uniform bandwidth. In modern ML clusters, the endpoint has shifted to individual GPUs, which are connected through a two-tier fabric consisting of faster intra-server (scale-up) links and slower inter-server (scale-out) links.111The terms ‘scale-up’ and ‘intra-server’ network are used interchangeably, and likewise ‘scale-out’ and ‘inter-server’. Consequently, All-to-All has become a key operation for many ML applications, including recommendation models [dlrm1, dlrm2], Gaussian Splatting [gaussiansplatting], and mixture-of-experts (MoE) models [deepspeed_moe, gshard, deepseek, tutel].
@@
The role of All-to-All in ML applications is important. In mixture-of-experts (MoE) models in particular, its cost can account for a large fraction of training time. MoE improves the efficiency of model parameters by activating only a subset of experts for each input token rather than all simultaneously. This selectivity, however, necessitates frequent All-to-All operations to dispatch tokens to experts across GPUs and aggregate their outputs. Prior studies [mixnet, tutel] show that MoE All-to-All can consume 30–55% of training time, making it a major contributor to overhead in large-scale training.
@@
The challenges of All-to-All arise at both the application and system layers. At the application layer, traffic is often skewed and dynamic. In MoE, some experts are selected more frequently than others, leading to larger data transfers for their corresponding GPUs during All-to-All. This imbalance keeps certain GPUs and NICs busy after others have finished, creating straggler effects. When communication volumes differ across endpoints in All-to-All, the operation is referred to as alltoallv [alltoallv]. Further, the traffic pattern in alltoallv changes every few hundred milliseconds, as the MoE gating function reassigns tokens to experts at runtime (Figure˜1). As a result, a GPU that is a hotspot at one moment may be idle the next. This dynamism makes static schedules impractical and requires schedulers to adapt in real time.
@@
At the system layer, the hardware fabric that connects GPUs is inherently two-tiered: fast intra-server links (scale-up) and much slower inter-server links (scale-out) (Figure˜4). This heterogeneity means that flows of the same size can finish quickly inside a server but take an order of magnitude longer across servers, leaving schedulers to coordinate thousands of flows over mismatched bandwidths. In addition, alltoallv’s dense communication pattern naturally triggers incast, a classic networking problem where many senders overload downlink of the same receiver. Incast causes network congestion with switch queue buildup and reduced goodput—even under modern congestion control—and remains an open challenge in large-scale cluster networking [ultraethernet-whitepaper, ultraethernet].
@@
Together, these challenges make efficient alltoallv scheduling especially difficult under the tight timescales demanded by MoE. Existing schedulers such as TACCL [taccl] and TE-CCL [teccl] employ solvers [gurobi] to generate near-optimal schedules. While they overcome the inefficiencies of fixed schedules in collective communication libraries such as NCCL and RCCL, their resulting formulations are NP-hard [taccl] and can take minutes to hours to synthesize a schedule even for just 32 GPUs. The state-of-the-art scheduler SyCCL [syccl] accelerates this process with heuristics and parallelism, yet requires seconds to minutes for balanced All-to-All, while skewed alltoallv still remains unresolved. Consequently, these schedulers are far too slow for MoE alltoallv workloads that shift every few hundred milliseconds.
@@
In this paper, rather than increasing scheduler complexity, we simplify the problem itself: It suffices to focus on optimizing the scale-out tier—the real bottleneck. We observe that scale-up is roughly an order-of-magnitude faster than scale-out (Figure˜4(b)). So the much faster scale-up fabric can cheaply absorb skew within each server, reshaping traffic before it reaches scale-out. With this, we can then maximize scale-out efficiency by keeping bottleneck servers transmitting at full rate while avoiding incast. Achieving this requires pairing senders and receivers without contention, which reduces naturally to a one-to-one matching problem between endpoints, solvable in polynomial time.
@@
Building on this insight, we propose FAST, a polynomial-time, matching-based scheduler for skewed and dynamic alltoallv workloads. FAST operates in two phases: (i) Skew mitigation, where it uses fast scale-up fabric to rebalance the scale-out workload so that all NICs face equal volume before traversing the slow scale-out links; and (ii) Balanced, one-to-one transfers, which use Birkhoff’s decomposition [birkhoff] to generate successive matchings between senders and receivers, ensuring scale-out transfers proceed without incast while keeping bottleneck servers fully active at line rate until completion. While Birkhoff’s decomposition has appeared in the design of switches [birkhoff_switch1, birkhoff_switch2, chronos], to our knowledge, this is the first work to apply Birkhoff’s decomposition to scheduling collective communication at GPU endpoints.
@@
We implement FAST on both NVIDIA H200 [h200] and AMD MI300X [mi300x] testbeds and evaluate it against state-of-the-art solutions such as DeepEP [deepep], TACCL [taccl], and TE-CCL [teccl]. Under skewed workloads, FAST outperforms the strongest NVIDIA baseline by 1.01–1.3 and the strongest AMD baselines by 1.5–2.8, and when integrated into Megatron-LM [megatron-lm] on AMD, improves end-to-end MoE training throughput by over RCCL [rccl] that suffers heavily from incast. The scheduler is highly efficient, completing in 221 s for 64 GPUs—fast enough for MoE workloads where the traffic matrix changes every few hundred milliseconds. Our code is available at [fast_github].
@@ 2 Motivation
All-to-All (alltoallv) communication—where every endpoint exchanges distinct data with all others—has become a key primitive in modern GPU clusters. In this setting, the communication endpoint is each individual GPU, connected by a two-tier fabric consisting of fast intra-server (scale-up) links and slower inter-server (scale-out) links.
@@
All-to-All communication cost in MoE models. Mixture-of-experts (MoE) is now a leading architecture for scalable large language models: instead of activating the full model for every input token, a lightweight gating network selects only a subset of ‘experts’, each implemented as a feed-forward network (FFN). While expert parallelism (EP) improves parameter efficiency, it introduces frequent, large-scale alltoallv operations to dispatch tokens to selected experts—often spanning hundreds of GPUs [perplexityai, deepseek]—and to gather results. As illustrated in Figure˜1, each MoE layer invokes alltoallv twice, and MoE layers often constitute a large fraction of the model. Prior measurements [tutel, mixnet] have shown that alltoallv can account for 30–56% of training time.
@@
While MoE is our primary focus, the importance of alltoallv extends beyond. It underpins recommendation systems [dlrm1, dlrm2], Gaussian Splatting [gaussiansplatting], and classical scientific workloads such as 3D FFT [fft] (where it can dominate up to 97.3% of runtime [fftoverhead]!). Despite being just one collective, alltoallv performance disproportionately shapes the efficiency of both modern AI and traditional HPC workloads.
@@
Application challenges—skewness & dynamism. To study this communication, we profile MoE training using Megatron-LM [megatron-lm] with 32 experts (one per GPU). We find that alltoallv workloads are inherently skewed and dynamic, consistent with recent profiling results of Mixtral models [mixnet, mixtral]. Unlike balanced collectives such as All-Reduce, MoE alltoallv generates a highly uneven demand matrix: some GPU pairs exchange more than 12 the median volume (Figure˜2(a)). This skew creates stragglers—NICs that remain busy long after others have finished, delaying the entire collective and stalling training progress (Figure˜3(a)). In this setting, scheduling is critical: by routing part of a heavy GPU’s traffic to idle NICs, a smart scheduler can smooth out skew and mitigate stragglers, as illustrated in Figure˜3(b).
@@
MoE workloads are also dynamic: alltoallv traffic pattern changes every few hundred milliseconds. As shown in Figure˜2(b), the traffic volume of a given GPU pair can vary significantly across successive alltoallv invocations, since token routing is jointly decided by the input tokens and per-MoE-layer gating functions (Figure˜1) and cannot be predicted in advance. Therefore, communication schedules must be recomputed online at the timescale as workload changes—a GPU pair with heavy load in one alltoallv may be nearly idle in the next—making fast, online scheduling essential.
@@
System challenges—heterogeneity & incast. Modern GPU clusters further complicate scheduling with two system-level challenges. First, a heterogeneous, two-tier fabric connects GPUs (Figure˜4(a)): fast intra-server links (e.g., 5th-gen NVLink 900 GBps [h100]) and much slower inter-server links (e.g., Ethernet 800 Gbps [cx8]). This two-tier network increases scheduling complexity: for each alltoallv, schedulers must navigate thousands of flows across mismatched bandwidths, exploring a large space of routing and pacing decisions, which can turn scheduling itself into a bottleneck.
@@
Second, incast is a classic networking problem that arises from alltoallv’s dense communication pattern, where many flows converge on the same NIC downlink in the scale-out fabric. While the burstiness of small messages can be absorbed by switch queues, MoE alltoallv transfers are much larger—typically 100 MB to 1 GB [deepep]—causing sustained congestion that requires active control. Even with advanced schemes [dcqcn, timely, ultraethernet], incast often results in unfair bandwidth sharing and degraded goodput. It remains an open challenge [ultraethernet-whitepaper], and schedulers often attempt to mitigate it proactively rather than relying on the transport layer alone.
@@
Limitations of existing approaches. State-of-the-art schedulers such as TACCL [taccl], TE-CCL [teccl], and SyCCL [syccl] are designed to be general: they support a wide range of collectives (All-Reduce, All-Gather, All-to-All, etc.) and reason about arbitrary topologies. To achieve this generality, they cast scheduling as constraint-satisfaction or optimization problems, often NP-hard [taccl], and solve them with heavy computation. This is practical for collectives with repetitive communication patterns like All-Reduce, where the high scheduling cost can be amortized over many iterations.
@@
For alltoallv, however, existing approaches are too slow. SyCCL [syccl], the fastest to date with parallelism and heuristic acceleration, is orders of magnitude faster than earlier solver-based systems but still requires minutes to produce a schedule for 64 GPUs—impractical when MoE traffic shifts every few hundred milliseconds. While such systems excel at achieving near-optimal completion under arbitrary topologies, their fine-grained modeling makes them hardly scalable to this setting.
@@
At the other extreme, production libraries like NCCL [nccl] generate schedules instantly, but rely on fixed schedules oblivious to the dynamic, skewed workload, often resulting in lower throughput than what the hardware could achieve.
@@
Goal. Can we design a fast, online scheduler for alltoallv that sustains high performance? Instead of targeting arbitrary collectives or topologies, we focus on a specialized solution for alltoallv on today’s two-tier GPU clusters—where skewness, dynamism, asymmetry, and incast constrain existing schedulers. FAST integrates with existing libraries: the runtime dispatches alltoallv to FAST and uses conventional algorithms for other collectives.
@@ 3 Design Overview
To generate a separate schedule for each alltoallv invocation in modern GPU clusters, we take a step back. Rather than enumerating complex constraints, we start by first solving alltoallv on a simplified single-tier network and then generalize the solution to today’s asymmetric two-tier fabrics.
Starting point: alltoallv in a single-tier network. On a single-tier, full-bisection network with uniform link bandwidth, the objective is to maximize communication efficiency by avoiding incast and congestion. This is achieved by ensuring that, at any instant, each sender communicates with exactly one receiver and each receiver accepts data from exactly one sender. As a result, communication can be organized into stages, where each stage realizes a one-to-one matching between senders and receivers, and successive stages collectively exchange data across all sender–receiver pairs.
@@
To solve this problem, we observe that Birkhoff’s decomposition [birkhoff]—introduced in 1946 as a mathematical theorem—can be reinterpreted as an optimal scheduling strategy for alltoallv. Formally, the theorem states that any traffic matrix222The theorem applies to scaled doubly stochastic matrices; arbitrary matrices can be adapted, as described later. can be expressed as a weighted sum of permutation matrices. Viewed through the scheduling lens, each permutation corresponds to a transfer stage: every active row (sender) and column (receiver) has exactly one nonzero entry of equal size (transfer size), so each participant exchanges data with exactly one partner and all finish the stage together. By summing over these permutation matrices, all flows advance in proportion to their demand, driving the transfer to completion.
Figure˜5 illustrates our strawman approach: the top pane shows the decomposition into (partial) permutation matrices, while the bottom pane shows the corresponding transfers. This approach is appealing because (i) completion time hits the lower bound—the bottleneck node (e.g., as sender in blue) transmits in all stages; (ii) each stage is balanced until nodes finish, so participants advance to the next stage together; and (iii) the decomposition is computationally efficient.
Challenges of applying Birkhoff’s decomposition to two-tier networks. Unfortunately, modern GPU clusters deviate from the simplified network setting in two ways. (i) In a two-tier fabric, even a ‘balanced’ permutation stage can finish unevenly—transfers on the inter-server links lag behind, idling the faster intra-server links and wasting bandwidth. (ii) In 8-GPU-per-server (e.g., HGX [hgx]) clusters, completion time is dictated by the busiest inter-server endpoints. Because most GPU pairs span server, the faster intra-server tier is rarely limiting. Under heavy skew, Birkhoff-based schedules cannot relieve this inter-server bottleneck and therefore still stall.
@@
Inter-server scheduling: balanced one-to-one transfers (§4.2). Once intra-server skew is absorbed, the remaining challenges are server-level imbalance and incast. Here, we apply Birkhoff’s decomposition to construct successive one-to-one, balanced transfer stages, ensuring bottleneck servers remain active at line rate until their traffic is complete.